OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y=4/9x-2 that passes through the point (4,3)?

OpenStudy (anonymous):

Oh this is easy.

OpenStudy (anonymous):

So, the slope of the equation is 4/9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You simply take the slope and

OpenStudy (anonymous):

Ugh, t\

OpenStudy (anonymous):

It wont'

OpenStudy (anonymous):

my backspace is broken , sorry

OpenStudy (anonymous):

Okay, so

OpenStudy (anonymous):

You take the slope which is : 4/9

OpenStudy (anonymous):

Take the negative reciprocal of it.

OpenStudy (anonymous):

Reciprocal means you sflip it.

OpenStudy (anonymous):

so if its 4/9 u flip it to 9/4 than make it a negative and it will be : -9/4

OpenStudy (anonymous):

so now you have -9/4 which is the slope

OpenStudy (anonymous):

now u use the equation y-yz

OpenStudy (anonymous):

UGH

OpenStudy (anonymous):

So - 9 / 4 x

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

you have the cordinates (4,3)

OpenStudy (anonymous):

x being 4 and y being 3

OpenStudy (anonymous):

so it will look like

OpenStudy (anonymous):

y-3=-9/4(x-4)

OpenStudy (anonymous):

distribute -9/4

OpenStudy (anonymous):

-9/4x and 28/4 and you add three to both sides

OpenStudy (anonymous):

which turns out to be

OpenStudy (anonymous):

y = -9/4x + 28/4 + 3

OpenStudy (anonymous):

28/4 is 7 so

OpenStudy (anonymous):

y = -9/4x + 7 + 3

OpenStudy (anonymous):

y = -9/4x + 10

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

But that isn't an answer choice on here...

OpenStudy (anonymous):

let me do it on my paper rq

OpenStudy (anonymous):

Okay thank you iv been stuck on this question for awhile.

OpenStudy (anonymous):

Ooops/

OpenStudy (anonymous):

Okay so -9/4 * -4 is 36/3 which turns out to be 12

OpenStudy (anonymous):

I accidentally said 28/4

OpenStudy (anonymous):

So we have y = -9/4x + 12 + 3

OpenStudy (anonymous):

y = -9/4x + 15

OpenStudy (anonymous):

Did you get that?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

lol np